Circular No. 6710

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1755-324 R. Remillard and A. Levine, Massachusetts Institute of Technology

(MIT); and J. Swank and T. Strohmayer, Goddard Space Flight Center (GSFC), report for the RXTE ASM Team at MIT and GSFC: “The All Sky Monitor of the Rossi X-ray Timing Explorer detected a new x-ray source late on July 25 UT, whose position was determined with the PCA early on July 29 to be α = 17^{h}55^{m}28^{s}.6, δ = −32^{o}28′39′′ (equinox 2000.0; uncertainty 1′). By July 28, the flux had increased to 170 mCrab (2–12 keV). The x-ray spectrum is substantially steeper than that of the Crab nebula. The spectrum appears complex, roughly consistent with a 0.7-keV disk blackbody combined with a hard component that extends above 10 keV. The flux is 2.9 × 10^{−9} erg cm^{−2} s^{−1} at 2–20 keV. The interstellar column density appears to be substanitally less than 1.0 × 10^{22} atoms cm^{−2}. Optical and radio observations are strongly encouraged.”

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD P. Garnavich, R. Kirshner, and P. Challis, Harvard-Smithsonian Cen-

ter for Astrophysics, report on behalf of the Supernova Intensive Study team that images of SN 1987A and its circumstellar emission rings were obtained with the Hubble Space Telescope (HST) Wide Field and Planetary Camera 2 (WFPC2) on July 10. A comparison between the narrow-band Hα (F656N) image and a WFPC2 image taken with the same filter in 1994 Feb. shows that a compact region near the inner ring has significantly increased in brightness (50 ± 5 percent), while the rest of the ring emission continues to fade slowly. The region is located 0′′.55 ± 0′′.02 from the supernova at p.a. 29^{o} ± 2^{o} and appears unresolved. This position is projected slightly within the ring, being 87 ± 4 percent of the separation between the supernova and the peak of the ring emission. The position of the bright spot is coincident with the blue-shifted-emission feature detected with STIS (IAUC 6665). Broad-band F675W (R) WFPC2 images show that the region remained at a nearly constant brightness between 1994 Oct. and 1996 Feb., while in the past 16 months the feature has brightened by 33 ± 5 percent (to R = 21.7). A similar increase in brightness is seen in the F336W (U), F439W (B), F555W (V), and F814W (I) filters, as well as the F502N filter (which isolates the [O III] 500.7-nm emission line). The sudden flux increase, combined with the spectral character and position of the spot within the ring, suggests that this is the first optical evidence of the collision between the supernova ejecta and the slow-moving progenitor wind.”

1997 July 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green